OpenStudy (anonymous):

(8 x -3)^2 x 13-6(14-10)^2 -1 explain.

OpenStudy (anonymous):

Are all of those x's multiplication? Please use asterisks if you're writing out multiplication because we use x a lot in algebra :)

OpenStudy (anonymous):

6b^2*(-b^7) so do i add the exponents?

OpenStudy (anonymous):

because it says to put the product in simplest form and the exponents are different so yeah do i add them ?

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

am i doing this right?

OpenStudy (anonymous):

which one are you doing? this one 6b^2*(-b^7)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so its a yes.. \((b^m)(b^n)=b^{m+n}\)

OpenStudy (anonymous):

oh so then its -6^9?

OpenStudy (anonymous):

Following from @Data_LG2's axiom, the negative just gets taken out the front (it may seem obvious but wanted to post it just in case!). \[(b^m)(-b^n)=-b^{m+n}\]

OpenStudy (anonymous):

\[(6b^2)(-b^7) = -6b^{7+2} = -6b^9\]

OpenStudy (anonymous):

i'm showing the general law of exponent but anyway, I agree with @tom982 .. thanks for that @mcnabkt don't forget the variable 'b'

OpenStudy (anonymous):

what about -7x^3y^5 *(-7y^3)

OpenStudy (anonymous):

What do you think the answer is @mcnabkt ? We've talked you through one already, so hopefully you'll have an idea of how to approach this question.

OpenStudy (anonymous):

well it takes a while for my brain to learn math like i cant look at big equations then learn it once and be an expert i have to have help through out the problem

OpenStudy (anonymous):